Dale invested his savings in two investment funds. The $8000 he invested in Fund A return a 10% profit. The amount he invested in Fund B returned a 3% profit. How much did he invest in Fund B, if both funds together returned a 7% profit?
Answer by richwmiller(17219)

.10*8000+.03*x=.07*(8000+x)
we know 3 rates and one amount invested
0.1*8000+0.03(x)=0.07(8000+x)
Multiply out.
800+0.03(x)=560+0.07x
combine like terms
-0.04x=-240
isolate x
x=-240/-0.04
Divide
x=6000
check
0.1*8000+0.03(x)=0.07(8000+x)
800+0.03*6000=0.07(8000+6000)
800+180=0.07(14000)
980=980
ok